# Enhancing the Cross-Size Generalization for Solving Vehicle Routing Problems via Continual Learning

### v1

@@ Abstract
Exploring machine learning techniques for addressing vehicle routing problems has attracted considerable research attention. To achieve decent and efficient solutions, existing deep models for vehicle routing problems are typically trained and evaluated using instances of a single size. This substantially limits their ability to generalize across different problem sizes and thus hampers their practical applicability. To address the issue, we propose a continual learning based framework that sequentially trains a deep model with instances of ascending problem sizes. Specifically, on the one hand, we design an inter-task regularization scheme to retain the knowledge acquired from smaller problem sizes in the model training on a larger size. On the other hand, we introduce an intra-task regularization scheme to consolidate the model by imitating the latest desirable behaviors during training on each size. Additionally, we exploit the experience replay to revisit instances of formerly trained sizes for mitigating the catastrophic forgetting. Experimental results show that our approach achieves predominantly superior performance across various problem sizes (either seen or unseen in the training), as compared to state-of-the-art deep models including the ones specialized for generalizability enhancement. Meanwhile, the ablation studies on the key designs manifest their synergistic effect in the proposed framework.
@@ 4 Methodology
Continual learning has emerged as a powerful approach for handling sequential tasks, which enables deep models to progressively retain and accumulate knowledge from evolving data streams. As illustrated in Figure 1, we harness CL to enhance the cross-size generalization capability of an autoregressive deep model $\theta$ (e.g., POMO (Kwon et al., 2020) (see Appendix E for details)), by sequentially training it on VRP instances of ascending problem sizes $\{N_{1},N_{2},...,N_{K}\}$. To ensure general favorable performance across the size spectrum, each size (i.e., task) $N_{i}$ (i=1,,Ki=1,...,K) is considered equally important and trained with the same task interval, which is defined as $E_{p}=E/K$ epochs where $E$ denotes the total training epochs of CL. In each task interval, the model is trained on each size to optimize the task-specific objective. Meanwhile, our approach exploits experience replay strategy to revisit instances of previously trained smaller sizes, so as to mitigate the catastrophic forgetting. Moreover, the inter-task or intra-task regularization scheme foster the model in current interval to emulate an exemplary model derived from previous or current interval, so as to inherit the previous learned knowledge. In this sense, our CL approach facilitates a coherent continuum of learning across varying problem sizes, which is elaborated in the following sections.
@@ 5.2 Generalization analysis
We also extend the evaluation to well-established benchmark datasets TSPLIB (Reinelt, 1991) and CVRPLIB (Uchoa et al., 2017) with size smaller than 500. The detailed results are reported in Table 6 and Table 7, respectively, where the gaps are calculated based on the optimal solution values for the instances. POMO-60, POMO-100, AMDKD-POMO, POMO-random, ELG-60 and ELG-100 are omitted due to their clear inferiority to POMO-150 and ELG-150 as shown in Table 1 and Table 2. Notably, benchmark datasets usually encompass a diverse range of sizes and customer distribution patterns. For our approach, we emphasize the intra-task regularization given its decent performance demonstrated earlier. The results show that Ours-POMO surpasses AMDKD-POMO that focuses on enhancing cross-size generalization. Although Omni-POMO is specially designed to enhance generalization across both size and distribution, our approach still delivers competitive performance to Omni-POMO. This further showcases the efficiency and effectiveness of our approach. Owing to meticulous designs tailored to bolster generalization on complex node distributions and large problem sizes, ELG-150 surpasses POMO-150 on both TSP and CVRP. Additionally, Ours-POMO and Ours-ELG outstrip their respective backbone baselines, i.e., POMO-150 and ELG-150, further demonstrating the effectiveness of our approach in enhancing cross-size generalizability (see Appendix B.1 for details).
@@
We further extend the evaluation to realistic data taken from “World TSP”, which is available at (rea, ) to show that both Ours-POMO and Ours-ELG consistently surpass their respective backbone baselines, i.e., POMO-150 and ELG-150, which further showcases the effectiveness of our approach (see Appendix B.2 for details).
@@ 5.3 Ablation study
In Table 4, we conduct an ablation study to clarify the effectiveness of each component of our approach on TSP and CVRP (refer to Appendix C.1), where only one regularization scheme can be used in our approach to keep a stable update of the exemplar model. The markers $\checkmark$” and $\times$” denote the utilization or exclusion of the corresponding component, respectively. The gaps are calculated based on the solutions acquired by Concorde in Table 1. As exhibited, experience replay, inter-task and intra-task regularization schemes contribute to the reduction of objective values and optimality gaps across all sizes, affirming their effectiveness in enhancing cross-size generalizability. Further combining them together, both Ours-inter and Ours-intra (last two rows) achieve better performance.
We further conduct an ablation study to demonstrate the effectiveness of the proposed regularization schemes, including both the inter-task regularization scheme and the intra-task one (see Appendix C.2 for details).
We also conduct experiments to show that our approach can consistently improve the performance across diverse sizes as the training progresses. To this end, we use the obtained models after training on each size to evaluate on both seen and unseen sizes. We present the details in Appendix D.

### v2

@@ Abstract
Deep models for vehicle routing problems are typically trained and evaluated using instances of a single size, which severely limits their ability to generalize across different problem sizes and thus hampers their practical applicability. To address the issue, we propose a continual learning based framework that sequentially trains a deep model with instances of ascending problem sizes. Specifically, on the one hand, we design an inter-task regularization scheme to retain the knowledge acquired from smaller problem sizes in the model training on a larger size. On the other hand, we introduce an intra-task regularization scheme to consolidate the model by imitating the latest desirable behaviors during training on each size. Additionally, we exploit the experience replay to revisit instances of formerly trained sizes for mitigating the catastrophic forgetting. Extensive experimental results show that the proposed approach achieves predominantly superior performance across various problem sizes (either seen or unseen in the training), as compared to state-of-the-art deep models including the ones specialized for the generalizability enhancement. Meanwhile, the ablation studies on the key designs manifest their synergistic effect in the proposed framework.
@@ 4 Methodology
Continual learning has emerged as a powerful approach for handling sequential tasks, which enables deep models to progressively retain and accumulate knowledge from evolving data streams. As illustrated in Figure 1, we harness CL to enhance the cross-size generalization capability of an autoregressive deep model $\theta$ (e.g., POMO (Kwon et al., 2020)), by sequentially training it on VRP instances of ascending problem sizes $\{N_{1},N_{2},...,N_{K}\}$. To ensure general favorable performance across the size spectrum, each size (i.e., task) $N_{i}$ (i=1,,Ki=1,...,K) is considered equally important and trained with the same task interval, which is defined as $E_{p}=E/K$ epochs where $E$ denotes the total training epochs of CL. In each task interval, the model is trained on each size to optimize the task-specific objective. Meanwhile, our approach exploits experience replay strategy to revisit instances of previously trained smaller sizes, so as to mitigate the catastrophic forgetting. Moreover, the inter-task or intra-task regularization scheme foster the model in current interval to emulate an exemplary model derived from previous or current interval, so as to inherit the previous learned knowledge. In this sense, our CL approach facilitates a coherent continuum of learning across varying problem sizes, which is elaborated in the following sections.
@@ 5.2 Generalization analysis
We further extend the evaluation to realistic data taken from “World TSP”, which is available at (rea, ) to show that both Ours-POMO and Ours-ELG consistently surpass their respective backbone baselines, i.e., POMO-150 and ELG-150, which further showcases the effectiveness of our approach.
@@ 5.3 Ablation study
In Table 4, we conduct an ablation study to clarify the effectiveness of each component of our approach on TSP, where only one regularization scheme can be used in our approach to keep a stable update of the exemplar model. The markers $\checkmark$” and $\times$” denote the utilization or exclusion of the corresponding component, respectively. The gaps are calculated based on the solutions acquired by Concorde in Table 1. As exhibited, experience replay, inter-task and intra-task regularization schemes contribute to the reduction of objective values and optimality gaps across all sizes, affirming their effectiveness in enhancing cross-size generalizability. Further combining them together, both Ours-inter and Ours-intra (last two rows) achieve better performance.